lindsaylynn1120:

guys another math problem Michelle has $12 and wants to buy a combination of dog food to feed at least three dogs at the animal shelter. A serving of dry food costs $1, and a serving of wet food costs $4. Part A: Write the system of inequalities that models this scenario. (5 points) Part B: Describe the graph of the system of inequalities, including shading and the types of lines graphed. Provide a description of the solution set. (5 points)

youngtringotringo:

lindsy lohan

lindsaylynn1120:

@youngtringotringo ????

youngtringotringo:

arieonna.

lindsaylynn1120:

@youngtringotringo still??? who are u????

Arieonna:

@alicole123 @laylalyssa can try to help u

lindsaylynn1120:

@arieonna wdym??

Arieonna:

try to mesage @laylalyssa and ask her if she can help i believe she is good at math

lindsaylynn1120:

@arieonna ohhh alright thxx

KyledaGreat:

Part A: Let x be the number of servings of dry food, and y be the number of servings of wet food. To feed at least three dogs, we have the inequality: x + y ≥ 3 To stay within the budget of $12, we have the inequality: x + 4y ≤ 12 So the system of inequalities that models this scenario is: x + y ≥ 3 x + 4y ≤ 12 Part B: To graph the system of inequalities, we can start by graphing the boundary lines for each inequality. For the first inequality, x + y ≥ 3, we can plot the line x + y = 3. This line has a y-intercept of 3 and a slope of -1, so we can plot two points on the line as (0, 3) and (3, 0), and draw a dashed line passing through these points. For the second inequality, x + 4y ≤ 12, we can plot the line x + 4y = 12. This line has a y-intercept of 3 and a slope of -1/4, so we can plot two points on the line as (0, 3) and (12, 0), and draw a dashed line passing through these points. Next, we need to shade the region that satisfies both inequalities. To do this, we can choose a test point that is not on either boundary line, such as (0, 0). If we substitute this point into each inequality, we get: (0) + (0) ≥ 3 (false) (0) + 4(0) ≤ 12 (true) Since the test point satisfies the second inequality but not the first, we need to shade the region that is below the line x + 4y = 12 and to the right of the line x + y = 3. We can shade this region and label it as the solution set. The lines are dashed because they represent inequalities, not equations. The solution set is shaded because it includes all the points that satisfy both inequalities. Therefore, the graph of the system of inequalities shows the feasible region for the number of servings of dry and wet food that Michelle can buy to feed at least three dogs within her budget of $12.

lindsaylynn1120:

@kyledagreat thank u so much

KyledaGreat:

You’re welcome